OpenStudy (anonymous):

what value of x satisfies the equation 8x + 4 = 220

OpenStudy (anonymous):

8x+4=220 -4 -4 8x=216 divide both sides by x = 27 8(27)+4=220? 216+4=220? 220=220

OpenStudy (anonymous):

*8x=216 divide both sides by 8